OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 2x − y = 4.

OpenStudy (lgbasallote):

x -intercept let y= 0 2x - 0 = 4 ^solve for x

OpenStudy (lgbasallote):

y-intercept let x = 0 2(0) - y = 4 0 - y =4 ^solve for y

OpenStudy (anonymous):

idk?